What are the [H3O+] and pH of 0.085M weak base C2H5NH2?



1
Expert's answer
2021-06-23T05:16:37-0400

"K_b=6.4\times10^{-4}"

"K_b=" "[C_2H_5NH_3^{+}][H_3O^+]\over[C_2H_5NH_2]"

"[0+x][0+x]\over[0.085-x]"

"K_b=" "[x][x]\over [0.085M]"

"6.4\times 10^{-4}=" "[x^2]\over [0.085]"

"6.4\times 10^{-4}[0.085]=x^2"

"\sqrt{(6.4\times 10^{-4}}[0.085]" "=\sqrt{x^2}"

"H_3O^{+}=0.00783M"

"pOH=-log[0.00738]"

"pOH=2.132"

"pH+pOH=14"

"pH=14-2.132=11.87"